OpenStudy (anonymous):

Which expression results from applying the Identity Property of Multiplication g+6g+3g+3h+4 to get an equivalent expression? A. g +g (6+3) +3h +4 B. g + 9g + 3h +4 C. g + 6g +3(g+h) +4 D. 7g +3g+3h+4

OpenStudy (anonymous):

I think its B.

OpenStudy (anonymous):

that's what I think too but that's just solving it ..,. im not sure if that is showing Identity Property of Multiplication

OpenStudy (anonymous):

Yea, I see what you are saying. I even tryed to do some research on Identity Property of Multiplication and all I really find is "The property that states the product of 1 and any number is that given number." I understand that. But this problem spins me in a whole different direction. But me personally, I would go with B. I'll try and get someone to help.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

haha yea same here and I ask Jim he said to give him 15 mins :)

OpenStudy (anonymous):

Oh ok. Well yea, see what he has to say. Sorry I wasn't able to be of much assistance hun.

jimthompson5910 (jim_thompson5910):

this problem is really vague, but my best guess is that you use the multiplicative property to go from g to 1g since 1*g = g then you use the distributive property to get 1g + 6g = (1+6)g = 7g

jimthompson5910 (jim_thompson5910):

so that would make the answer D